Exercise 2-7

Suppose the following information is available for Callaway Golf Company for the years 2017 and 2016. (Dollars are in thousands, except share information.)

2017 2016
Net sales $ 1,120,000 $ 1,126,100
Net income (loss) 65,300 53,600
Total assets 855,338 838,078
Share information
Shares outstanding at year-end 64,507,000 66,282,000
Preferred dividends 0 0


There were 73,139,000 shares outstanding at the end of 2015.

(a)

What was the company’s earnings per share for each year? (Round answers to 2 decimal places, e.g. 15.25.)

2017 : $

Entry field with incorrect answer

2016 : $

Entry field with incorrect answer
Answer #1

Answers

  • Requirement [a]

2017 :

$ 1.00

2016 :

$ 0.77

  • Workings

Working #1 2016

A

2015 Shares outstanding

               73,139,000

B

2016 shares outstanding

               66,282,000

C = A + B

Total

            139,421,000

D = C/2

Weighted average share outstanding

               69,710,500

E

Net Income 2016

$53,600,000

F = E/D

Earnings per share

$0.77

Working #1 2017

A

2016 shares outstanding

                                66,282,000

B

2017 shares outstanding

                                64,507,000

C = A + B

Total

                             130,789,000

D = C/2

Weighted average share outstanding

                                65,394,500

E

Net Income 2017

$65,300,000

F = E/D

Earnings per share

$1.00